OpenStudy (anonymous):

How would I compute the sum for Ʃ ln (1/4) from 0 to inf

OpenStudy (anonymous):

ln(1/4) is a constant term, i think you might be missing something in this question

OpenStudy (anonymous):

oh wait i wrote the question wrong @Stom, I had the telescopic series: ln 1/2 + 2/3 + ln 3/4 and using the rule ln(a) + ln(b) = ln(ab) and got the value ln = 1/4 is there any furthur I should go with the question.

OpenStudy (anonymous):

what is the complete question?

OpenStudy (anonymous):

compute the sum for other telescoping series : ln 1/2 + ln 2/3 + 3/4